Softee Blankets charges $6 for shipping orders of $20 or less, $10 for orders from $20.01 to $70, and $15 for orders over $70. One week shipping charges for 540 orders totaled $5320. There were 50 more orders for $20 or less shipped than there were orders for over $70. Find the number of orders shipped at each rate.
So far I've got
x+y+z=31
50x-z=0
6x+10y+15z=?
I'm just having problems finding the third system. The result always comes out negative and you can't sell a negative amount of items so somewhere I messed up I just don't know where.

Answer by rwm(914)

Where did the 31 come from? Oh I see you added the dollar amounts of the charges.6,10,15-
Why are you multiplying 50 times x? It says 50 more not 50 times. Plus it is x that is more than z so you have to add to z to get it to equal x.
At least you tried.
Here you go.
x+y+z=540
x=z+50
6x+10y+15z=5320
Solve these equations to get your three values.
x=170
y=250
z=120